OpenStudy (sillybutt100):

Can somebody please help me figure out how to simplify radical expressions like this: (n+2)/(2n+4) I don't want the answer. I want to know how to solve it.

OpenStudy (anonymous):

The method for multiplying polynomial equations like this one will FOIL. Foil stands for: First, Outer, Inner, Last First you multiply the first values of the two polynomials: n*2n = 2n^2 Then you multiply the outer values: 4*n = 4n Then the inner values: 2*2n = 4n Then the last values: 2*4 = 8 And finally, you add them all together: 2n^2 + 4n + 4n + 8 or 2n^2 +8n + 8 I hope you will understand this explanation, but just in case, this website explains the method pretty well: http://www.algebrahelp.com/lessons/simplifying/foilmethod/

OpenStudy (sillybutt100):

But that isn't quite what i meant I accidentally put multiply instead of simplify the equation looks like this \[\frac{ n+2 }{ 2n+4 }\]

OpenStudy (noelgreco):

Factor out the ZGCF in the denominator.

OpenStudy (noelgreco):

GCF

OpenStudy (anonymous):

Oh. In that case you just try to simplify both the top and bottom as much as possible and then cancel out variables. \[\frac{ n +2 }{ 2(n+2)? }\] and you can cancel out the n+2 on from the numerator and denominator to get 1/2.

OpenStudy (sillybutt100):

Thank you!

OpenStudy (noelgreco):

Gee, you're welcome.

OpenStudy (sillybutt100):

One more question wouldyou solve this the same way? \[\frac{ 15x ^{^{2}}+21x }{ 12x ^{3} }\]

OpenStudy (anonymous):

Yes - just do as much simplifying as possible. Like Noel said, you are looking for the greatest common factor. The numerator can be factored into 3x(5x+7) And this 3x, is the greatest common factor between the numerator and the denominator. Therefore, you can cancel it out from both of them to get the simplified answer. = \[\frac{ 5x + 7 }{ 4x ^{2} }\]

OpenStudy (sillybutt100):

Thank you

OpenStudy (sillybutt100):

One more question I promise this is the last one. How do you do this one: \[\frac{ a ^{3} +3a-28}{a ^{2}-49 }\]